# Also at ]Alpine Quantum Technologies GmbH, Innsbruck, Austria

### v1

@@ Appendix A Stochastic processes and Monte Carlo methods
$\rho(t)=\mathbb{E}\big{[}\ket{\tilde{\psi}(t)}\!\!\bra{\tilde{\psi}(t)}\big{]}\approx\sum_{i,j}\left(\frac{1}{M_{\rm MC}}\sum_{n=1}^{M_{\rm MC}}\tilde{c}_{i,n}(t)\tilde{c}_{j,n}^{*}(t)\right)\ket{i}\!\!\bra{j},$ (64)
@@ Appendix B Time-convolutionless master equations and the cumulant expansion
such that the microscopic coupling strength is set by $\alpha=\Omega$. As discussed in [103, 141], the so-called cluster property of the $n$-th contribution to the kernel $\mathcal{K}_{n}(t)$ is crucial to identify a small truncation parameter. To understand it, let us start by focusing on $\mathcal{K}_{2}(t)$, which depends on the two-time auto-correlation functions in $\mathbb{E}[\tilde{\mathcal{L}}_{t}\tilde{\mathcal{L}}_{t^{\prime}}]$. We can thus identify $\hat{\tau}_{\rm c}$ as a correlation time such that, for times $|t-t^{\prime}|\gg\hat{\tau}_{\rm c}$, the auto-correlation functions in $\mathcal{K}_{2}(t)$ would become negligibly small. Here, a hat is used to distinguish this correlation time from that of the previous Appendix, $\tau_{\rm c}$. This correlation time would be used to model a colored dephasing noise $\delta\tilde{\omega}(t)$ on its own, regardless of its effect on the qubit. For the phase diffusion model mentioned in the main text [86, 87, 88, 89], a type of phase noise that can be traced back to the Anderson-Weiss model of random-frequency modulations [142] and to Kubo’s oscillator [143], the laser frequency fluctuates as a white noise process with diffusion constant $c$, such that the detuning in Eq (84) reads $\delta\tilde{\omega}(t)=\sqrt{c}\tilde{\xi}(t)$. This noise has a vanishing correlation time $\tau_{\rm c}=0$ and, following Eq. (57), a flat power spectral density $S(\omega)=c$. Moreover, its time integral is the Wiener random process $W(t)$ that directly affects the above auto-correlation function $\mathbb{E}[\tilde{\mathcal{L}}_{t}\tilde{\mathcal{L}}_{t^{\prime}}]$. In fact, using (57), we find an exponential law $\mathbb{E}[{\rm e}^{{\rm i}(W(t)-W(t^{\prime}))}]={\rm e}^{-\frac{c}{2}|t-t^{\prime}|}$ controlled by the noise diffusion constant. This yields the correlation time $\hat{\tau}_{\rm c}=2/c$, which clearly differs from that of the white frequency noise. In fact, it can be connected to the non-zero linewidth of the laser radiation which, within this model, has a Lorentzian PSD $S_{\rm L}(\omega)\propto E_{0}^{2}\int_{-\infty}^{\infty}{\rm d}s\,{\rm e}^{({\rm i}(\omega_{\rm L}-\omega)-\frac{c}{2})s}\propto cE_{0}^{2}/((\omega-\omega_{\rm L})^{2}+(c/2)^{2})$ that clearly differs from the flat PSD of the dephasing noise $S(\omega)$. The laser linewidth $\Delta\omega$, defined as the full width at half maximum (FWHM) of the PSD, is fixed by the diffusion constant $\Delta\omega=c=2/\hat{\tau}_{\rm c}$, such that $\hat{\tau}_{\rm c}$ can be interpreted as the finite correlation time of a quasi-monochromatic laser [89].
@@
where the effective dephasing time $T_{2}=2/S(0)$ depends on the value of the noise PSD evaluated at a vanishing frequency, which is consistent with the fact that the system does not absorb energy from the stochastic process. The condition of validity for the second-order truncation dressed-state master equation (12) is thus $\zeta_{\rm dress}\ll 1$ as higher-order terms will scale with even smaller $\zeta_{\rm dress}^{n}$, which hence requires setting $\tau_{\rm c}\ll T_{2}$ and no longer imposes any condition on the Rabi frequency and the gate duration. In this way, we can explore the regime of high-fidelity gates and coherent Rabi flops exploring timescales $t~{}\sim\frac{1}{\Omega}\gtrsim\tau_{\rm c}$. The important difference is that the effect of the dephasing can still be small for these timescale $t\ll T_{2}$, which is consistent with the regime of validity of our approximation $\zeta_{\rm dress}\ll 1$. We can thus describe single-qubit gates with a high fidelity even in fast regimes where the time-correlation effects in the noise are no longer negligible.
@@ Appendix C Solution of the dressed-state master equation
In this appendix, we give a detailed account for the analytic solution of Eq. (12). For $\phi=0$, the dressed-state basis is composed of |±$\ket{\pm}=\frac{1}{2}(\ket{0}\pm\ket{1})$, and the action of the unitary operator that changes the frame on the dressed-state basis states is $U_{\Omega}(t)\ket{\pm}={\rm exp}\{\mp{\rm i}\frac{\Omega}{2}t\}\ket{\pm}$. Using these expressions one can rewrite Eq. (12) in the dressed-state basis. After some algebra, the evolution for the populations, which correspond to the coherences in the original basis, satisfies
@@
$\gamma_{1}(t)=-\int_{0}^{t}\text{d}t^{\prime}C(t-t^{\prime})\cos(\Omega(t-t^{\prime})),$ (91)
@@
$\begin{split}\gamma_{2}(t)=-\int_{0}^{t}\!\!\text{d}t^{\prime}\,C(t-t^{\prime})\cos(\Omega(t+t^{\prime})),\\ \delta_{1}(t)=-\int_{0}^{t}\!\!\text{d}t^{\prime}\,C(t-t^{\prime})\sin(\Omega(t-t^{\prime})),\\ \delta_{2}(t)=+\int_{0}^{t}\!\!\text{d}t^{\prime}\,C(t-t^{\prime})\sin(\Omega(t+t^{\prime})).\end{split}$ (94)
@@ Appendix D Process error matrix and non-Markovianity
In this Appendix, we present the specific expressions for the error process matrix $\chi^{\rm err}(t)$ obtained from our analytical treatment of the dressed-state master equation of Sec. II.1. We recall that this master equation is obtained from the second-order cumulant expansion of the time-convolutionless kernel presented in B, when one works in an instantaneous frame that rotates with the Rabi frequency. In the main text, we discussed the QPT results for the Kraus operators, which have a simple expression (32) when restricting to a pair of filter functions leading to $\Gamma_{1}(t)$, and $\Delta_{1}(t)$ in Eq. (18). On the other hand, our full solutions also depend on the additional pair of filter functions $\Gamma_{2}(t)$ and $\Delta_{2}(t)$ in Eq. (18), which codify additional effects of the statistical properties of the noise given by $C(t-t^{\prime})$ via the filtered PSD. In this Appendix, we present the explicit expressions of the error process matrix taking into account these contributions and, furthermore, show that these can lead to a non-zero measure of non-Markovianity [108, 109]. It should be pointed out that in the quantum case, different definitions of non-Markovianity exist in the literature that –while not the present model– may not agree upon whether non-Markovianity is present in certain open-system dynamics.
@@
$\begin{split}\Gamma_{1}(t)&=\frac{1}{2}{S(\Omega)}\bigg{(}t-\tau_{\rm c}\frac{2(\Omega\tau_{\rm c})}{1+(\Omega\tau_{\rm c})^{2}}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\sin\Omega t-\tau_{\rm c}\frac{1-(\Omega\tau_{\rm c})^{2}}{1+(\Omega\tau_{\rm c})^{2}}(1-{\rm e}^{-\frac{t}{\tau_{\rm c}}}\cos\Omega t)\bigg{)},\\ \Gamma_{2}(t)&=\frac{1}{2}S(\Omega)\cos\Omega t\left(\frac{1}{\Omega}\sin\Omega t-\tau_{\rm c}\cos\Omega t+\tau_{\rm c}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\right),\\ \Delta_{1}(t)&=\frac{1}{2}S(\Omega)\bigg{(}t(\Omega\tau_{\rm c})+\tau_{\rm c}\frac{1-(\Omega\tau_{\rm c})^{2}}{1+(\Omega\tau_{\rm c})^{2}}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\sin\Omega t-\tau_{\rm c}\frac{2(\Omega\tau_{\rm c})}{1+(\Omega\tau_{\rm c})^{2}}(1-{\rm e}^{-\frac{t}{\tau_{\rm c}}}\cos\Omega t)\bigg{)},\\ \Delta_{2}(t)&=\frac{1}{2}S(\Omega)\left(\frac{1}{\Omega}\sin^{2}\Omega t-\frac{\tau_{\rm c}}{2}\sin 2\Omega t+\tau_{\rm c}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\right).\end{split}$ (106)
@@
$\mathcal{D}_{t}(\rho)=\sum_{\alpha}\bar{\gamma}_{\alpha}({t^{\prime}})\!\!\left(\!L_{\alpha}({t})\rho L^{\dagger}_{\alpha}({t})-\textstyle\frac{1}{2}\{L_{\alpha}(t)L^{\dagger}_{\alpha}(t),\rho\}\!\!\right)\!\!.$ (112)
@@
This measure is depicted in Fig. 3, where we see how it increases with time for all the Rabi frequencies until it saturates at a given time $t^{*}$, which is the same one for all the frequencies. It shows that non-Markovianities are bounded to a time region whose length is modulated by the correlation time $\tau_{c}$ of the noise process. It is also remarkable how the importance of the non-Markovian effects grows with the Rabi frequency until a maximum value around $\Omega=5\rm{rad/s}$, and then decreases again for higher values of the frequency.

### v3

@@ Appendix A Stochastic processes and Monte Carlo methods
$\rho(t)=\mathbb{E}\big{[}\ket{\tilde{\psi}(t)}\!\!\bra{\tilde{\psi}(t)}\big{]}\approx\sum_{i,j}\left(\frac{1}{M_{\rm MC}}\sum_{n=1}^{M_{\rm MC}}\tilde{c}_{i,n}^{\phantom{*}}(t)\tilde{c}_{j,n}^{*}(t)\right)\ket{i}\!\!\bra{j},$ (77)
@@ Appendix B Time-convolutionless master equations and the cumulant expansion
such that the microscopic coupling strength is set by $\alpha=\Omega$. As discussed in [103, 142], the so-called cluster property of the $n$-th contribution to the kernel $\mathcal{K}_{n}(t)$ is crucial to identify a small truncation parameter. To understand it, let us start by focusing on $\mathcal{K}_{2}(t)$, which depends on the two-time auto-correlation functions in $\mathbb{E}[\tilde{\mathcal{L}}_{t}\tilde{\mathcal{L}}_{t^{\prime}}]$. We can thus identify $\hat{\tau}_{\rm c}$ as a correlation time such that, for times $|t-t^{\prime}|\gg\hat{\tau}_{\rm c}$, the auto-correlation functions in $\mathcal{K}_{2}(t)$ would become negligibly small. Here, a hat is used to distinguish this correlation time from that of the previous Appendix, $\tau_{\rm c}$. This correlation time would be used to model a colored dephasing noise $\delta\tilde{\omega}(t)$ on its own, regardless of its effect on the qubit. For the phase diffusion model mentioned in the main text [86, 87, 88, 89], a type of phase noise that can be traced back to the Anderson-Weiss model of random-frequency modulations [143] and to Kubo’s oscillator [144], the laser frequency fluctuates as a white noise process with diffusion constant $c$, such that the detuning in Eq (97) reads $\delta\tilde{\omega}(t)=\sqrt{c}\tilde{\xi}(t)$. This noise has a vanishing correlation time $\tau_{\rm c}=0$ and, following Eq. (70), a flat power spectral density $S_{\tilde{\omega}}(\omega)=c$. Moreover, its time integral is the Wiener random process $W(t)$ that directly affects the above auto-correlation function $\mathbb{E}[\tilde{\mathcal{L}}_{t}\tilde{\mathcal{L}}_{t^{\prime}}]$. In fact, using (70), we find an exponential law $\mathbb{E}[{\rm e}^{{\rm i}(W(t)-W(t^{\prime}))}]={\rm e}^{-\frac{c}{2}|t-t^{\prime}|}$ controlled by the noise diffusion constant. This yields the correlation time $\hat{\tau}_{\rm c}=2/c$, which clearly differs from that of the white frequency noise. In fact, it can be connected to the non-zero linewidth of the laser radiation which, within this model, has a Lorentzian PSD $S_{\rm L}(\omega)\propto E_{0}^{2}\int_{-\infty}^{\infty}{\rm d}s\,{\rm e}^{({\rm i}(\omega_{\rm L}-\omega)-\frac{c}{2})s}\propto cE_{0}^{2}/((\omega-\omega_{\rm L})^{2}+(c/2)^{2})$ that clearly differs from the flat PSD of the dephasing noise $S_{\tilde{\omega}}(\omega)$. The laser linewidth $\Delta\omega$, defined as the full width at half maximum (FWHM) of the PSD, is fixed by the diffusion constant $\Delta\omega=c=2/\hat{\tau}_{\rm c}$, such that $\hat{\tau}_{\rm c}$ can be interpreted as the finite correlation time of a quasi-monochromatic laser [89].
@@
where the effective dephasing time $T_{2}=2/S_{\tilde{\omega}}(0)$ depends on the value of the noise PSD evaluated at a vanishing frequency, which is consistent with the fact that the system does not absorb energy from the stochastic process. The condition of validity for the second-order truncation dressed-state master equation (12) is thus $\zeta_{\rm dress}\ll 1$ as higher-order terms will scale with even smaller $\zeta_{\rm dress}^{n}$, which hence requires setting $\tau_{\rm c}\ll T_{2}$ and no longer imposes any condition on the Rabi frequency and the gate duration. In this way, we can explore the regime of high-fidelity gates and coherent Rabi flops exploring timescales $t~{}\sim\frac{1}{\Omega}\gtrsim\tau_{\rm c}$. The important difference is that the effect of the dephasing can still be small for these timescale $t\ll T_{2}$, which is consistent with the regime of validity of our approximation $\zeta_{\rm dress}\ll 1$. We can thus describe single-qubit gates with a high fidelity even in fast regimes where the time-correlation effects in the noise are no longer negligible.
@@ Appendix C Solution of the dressed-state master equation
In this appendix, we give a detailed account for the analytic solution of Eq. (12) and Eq. (44). For $\phi=0$, the dressed-state basis is composed of |±$\ket{\pm}=\frac{1}{2}(\ket{0}\pm\ket{1})$, and the action of the unitary operator that changes the frame on the dressed-state basis states is $U_{\Omega}(t)\ket{\pm}={\rm exp}\{\mp{\rm i}\frac{\Omega}{2}t\}\ket{\pm}$. Using these expressions one can rewrite Eq. (12) in the dressed-state basis. After some algebra, the evolution for the populations, which correspond to the coherences in the original basis, satisfies
@@
$\gamma_{1}(t)=-\int_{0}^{t}\text{d}t^{\prime}C_{\tilde{\omega}}(t-t^{\prime})\cos(\Omega(t-t^{\prime})),$ (104)
@@
$\begin{split}\gamma_{2}(t)=-\int_{0}^{t}\!\!\text{d}t^{\prime}\,C_{\tilde{\omega}}(t-t^{\prime})\cos(\Omega(t+t^{\prime})),\\ \delta_{1}(t)=-\int_{0}^{t}\!\!\text{d}t^{\prime}\,C_{\tilde{\omega}}(t-t^{\prime})\sin(\Omega(t-t^{\prime})),\\ \delta_{2}(t)=+\int_{0}^{t}\!\!\text{d}t^{\prime}\,C_{\tilde{\omega}}(t-t^{\prime})\sin(\Omega(t+t^{\prime})).\end{split}$ (107)
@@
For the case with amplitude fluctuations, we follow exactly the same procedure. The new contribution enters solely in the non-diagonal terms:
$\begin{split}\frac{{\rm d}{\rho}_{+-}}{{\rm d}t}=-\frac{1}{2}\left[(\tilde{\gamma}_{1}(t)+{\rm i}\delta_{1}(t))\rho_{+-}-(\gamma_{2}(t)+{\rm i}\delta_{2}(t))\rho_{-+}\right],\\ \frac{{\rm d}{\rho}_{-+}}{{\rm d}t}=-\frac{1}{2}\left[(\tilde{\gamma}_{1}(t)-{\rm i}\delta_{1}(t))\rho_{+-}-(\gamma_{2}(t)-{\rm i}\delta_{2}(t))\rho_{-+}\right],\end{split}$ (112)
where $\delta_{1,2}$ are functions defined in Eq. (107) identical to the purely phase noise case, and the decay rate $\tilde{\gamma}_{1}(t)=\gamma_{1}(t)-\delta\gamma_{1}(t)$ has been redressed by an additional function $\delta\gamma_{1}$ given by
$\delta\gamma_{1}(t)=2\int_{0}^{t}{\rm{d}}t^{\prime}\,C_{\Omega}(t-t^{\prime})$ (113)
which acts as an induced decay rate. Hence the new system of equations reads as Eq. (108) but with the redressed decay rate $\tilde{\gamma}_{1}(t)$. The solution for the evolution of the coherences is then the same as in the pure dephasing noise but adding the new function
$\Delta\Gamma_{1}(t)=\int_{0}^{t}{\rm d}t^{\prime}\delta\gamma_{1}(t^{\prime}),$ (114)
which can be rewritten as in Eq. (46) of the main text after performing a cosine Fourier transform.
@@ Appendix D Process error matrix and non-Markovianity
In this Appendix, we present the specific expressions for the error process matrix $\chi^{\rm err}(t)$ obtained from our analytical treatment of the dressed-state master equation of Sec. II.1. We recall that this master equation is obtained from the second-order cumulant expansion of the time-convolutionless kernel presented in B, when one works in an instantaneous frame that rotates with the Rabi frequency. In the main text, we discussed the QPT results for the Kraus operators, which have a simple expression (32) when restricting to a pair of filter functions leading to $\Gamma_{1}(t)$, and $\Delta_{1}(t)$ in Eq. (18). On the other hand, our full solutions also depend on the additional pair of filter functions $\Gamma_{2}(t)$ and $\Delta_{2}(t)$ in Eq. (18), which codify additional effects of the statistical properties of the noise given by $C_{\tilde{\omega}}(t-t^{\prime})$ via the filtered PSD. In this Appendix, we present the explicit expressions of the error process matrix taking into account these contributions and, furthermore, show that these can lead to a non-zero measure of non-Markovianity [108, 109]. It should be pointed out that in the quantum case, different definitions of non-Markovianity exist in the literature that –while not the present model– may not agree upon whether non-Markovianity is present in certain open-system dynamics.
@@
Let us now discuss how the process matrix changes when including amplitude fluctuations. When considering multi-axis noise, the process matrix still retains a simple form as a block-diagonal matrix in the Pauli basis
$\tilde{\chi}^{\rm err}(t)=\frac{1}{4}\left(\begin{array}[]{c | c}\tilde{\chi}^{\rm err}_{A}(t)&\begin{array}[]{cc}0&0\\ 0&0\end{array}\\ \hline\cr\begin{array}[]{cc}0&0\\ 0&0\end{array}&\tilde{\chi}^{\rm err}_{B}(t)\end{array}\right).$ (119)
Here, the upper block can be expressed as
$\begin{split}\tilde{\chi}^{\rm err}_{A}(t)=&\left(1+{\rm e}^{-\Gamma_{1}(t)}\right)\mathbb{1}_{2}+2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\cos\textstyle\frac{1}{2}\Theta(t)\sigma_{z}\\ &-2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\frac{\Delta_{1}(t)}{\Theta(t)}\sin\textstyle\frac{1}{2}\Theta(t)\sigma_{y},\end{split}$ (120)
whereas the lower block of reads
$\begin{split}\tilde{\chi}^{\rm err}_{B}(t)=&\left(1-{\rm e}^{-\Gamma_{1}(t)}\right)\mathbb{1}_{2}+2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\frac{\Gamma_{2}(t)}{\Theta(t)}\sin\textstyle\frac{1}{2}\Theta(t)\sigma_{z}\\ &+2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\frac{\Delta_{2}(t)}{\Theta(t)}\sin\textstyle\frac{1}{2}\Theta(t)\sigma_{y}.\end{split}$ (121)
@@
$\begin{split}\Gamma_{1}(t)&=\frac{1}{2}{S_{\tilde{\omega}}(\Omega)}\bigg{(}t-\tau_{\rm c}\frac{2(\Omega\tau_{\rm c})}{1+(\Omega\tau_{\rm c})^{2}}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\sin\Omega t-\tau_{\rm c}\frac{1-(\Omega\tau_{\rm c})^{2}}{1+(\Omega\tau_{\rm c})^{2}}(1-{\rm e}^{-\frac{t}{\tau_{\rm c}}}\cos\Omega t)\bigg{)},\\ \Gamma_{2}(t)&=\frac{1}{2}S_{\tilde{\omega}}(\Omega)\cos\Omega t\left(\frac{1}{\Omega}\sin\Omega t-\tau_{\rm c}\cos\Omega t+\tau_{\rm c}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\right),\\ \Delta_{1}(t)&=\frac{1}{2}S_{\tilde{\omega}}(\Omega)\bigg{(}t(\Omega\tau_{\rm c})+\tau_{\rm c}\frac{1-(\Omega\tau_{\rm c})^{2}}{1+(\Omega\tau_{\rm c})^{2}}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\sin\Omega t-\tau_{\rm c}\frac{2(\Omega\tau_{\rm c})}{1+(\Omega\tau_{\rm c})^{2}}(1-{\rm e}^{-\frac{t}{\tau_{\rm c}}}\cos\Omega t)\bigg{)},\\ \Delta_{2}(t)&=\frac{1}{2}S_{\tilde{\omega}}(\Omega)\left(\frac{1}{\Omega}\sin^{2}\Omega t-\frac{\tau_{\rm c}}{2}\sin 2\Omega t+\tau_{\rm c}{\rm e}^{-\frac{t}{\tau_{\rm c}}}\right).\end{split}$ (125)
@@
$\mathcal{D}_{t}(\rho)=\sum_{\alpha}\bar{\gamma}_{\alpha}({t^{\prime}})\!\!\left(\!L_{\alpha}({t})\rho L^{\dagger}_{\alpha}({t})-\textstyle\frac{1}{2}\{L^{\phantom{\dagger}}_{\alpha}(t)L^{\dagger}_{\alpha}(t),\rho\}\!\!\right)\!\!.$ (131)
@@
This measure is depicted in Fig. 3 b, where we see how it increases with time for all the Rabi frequencies until it saturates at a given time $t^{*}$, which is the same one for all the frequencies. It shows that non-Markovianities are bounded to a time region whose length is modulated by the correlation time $\tau_{c}$ of the noise process. It is also remarkable how the importance of the non-Markovian effects grows with the Rabi frequency until a maximum value around $\Omega=5\rm{rad/s}$, and then decreases again for higher values of the frequency.
We can also study the non-Markovianity of the error channel when amplitude fluctuations are taken into account. Starting from the dressed-state master equation for the evolution of the populations and coherences, Eq. (44), one can again rewrite the evolution equation in the Pauli basis as
$d(t)=\textstyle\frac{1}{2}\begin{pmatrix}0&0&0\\ 0&\tilde{\gamma_{1}}(t)+\gamma_{2}(t)&-\delta_{2}(t)\\ 0&-\delta_{2}(t)&\tilde{\gamma_{1}}(t)-\gamma_{2}(t)\end{pmatrix},$ (138)
with the new dressed decaying rate $\tilde{\gamma_{1}}(t)=\gamma_{1}(t)+\delta\gamma_{1}(t)$.
By diagonalizing the above orthogonal matrix, one readily finds that the canonical time-dependent rates are
$\bar{\gamma}_{\alpha}(t)\in\left\{0,\textstyle\frac{1}{2}\tilde{\gamma_{1}}(t)\pm\textstyle\frac{1}{2}\sqrt{\gamma_{2}^{2}(t)+\delta_{2}^{2}(t)})\right\}.$ (139)
This readily generalizes our previous results via a renormalization of the purely dephasing noise decaying rate.